### v1

# PVPO: Pre-Estimated Value-Based Policy Optimization for Agentic Reasoning

Abstract
Critic-free reinforcement learning methods, particularly group policies, have attracted considerable attention for their efficiency in complex tasks. However, these methods rely heavily on multiple sampling and comparisons within the policy to estimate advantage, which may cause the policy to fall into local optimum and increase computational cost. To address these issues, we propose PVPO, an efficient reinforcement learning method enhanced by an advantage reference anchor and data pre-sampling. Specifically, we use the reference model to rollout in advance and employ the calculated reward score as a reference anchor. Our approach effectively corrects the cumulative bias introduced by intra-group comparisons and significantly reduces reliance on the number of rollouts. Meanwhile, the reference model can assess sample difficulty during data pre-sampling, enabling effective selection of high-gain data to improve training efficiency. Experiments conducted on nine datasets across two domains demonstrate that PVPO achieves State-Of-The-Art (SOTA) performance. Our approach not only demonstrates robust generalization across multiple tasks, but also exhibits scalable performance across models of varying scales.
1 Introduction
Reinforcement Learning (RL) is a machine learning method for learning optimal policies through interaction with the environment. Policy optimization depends on accurately estimating the advantage function to improve the agent’s actions. In classic actor-critic frameworks, a critic network predicts state-value ($V$), which combines with action-value ($Q$) to compute the advantage and then guides policy updates. Recently, research has increasingly focused on more efficient critic-free architectures. These methods do not directly compute the absolute advantage. Instead, they build baselines for relative advantage, simplifying the training process and reducing resource consumption (Shao et al., 2024; Feng et al., 2025b).
Grouping policies, as used in critic-free RL methods like GRPO (Shao et al., 2024), become an important research topic. This is not only because they demonstrate superior performance, but also because the removal of the value model saves training resources, enabling researchers to train larger-scale models under limited hardware conditions. Although PPO and other actor-critic methods sometimes achieve higher accuracy, critic-free grouping policies are widely used for their practical efficiency. Some studies group by sample, running multiple trajectories within each group to compute relative advantage (Zuo et al., 2025; Lyu et al., 2025). Others group by action or timestep, enabling finer partitioning and more accurate baseline estimation (Feng et al., 2025b; Li et al., 2025a). These methods can improve baseline accuracy for similar trajectories. However, grouping policies usually require more rollouts to boost performance, which greatly increases computational cost. Methods such as DAPO (Yu et al., 2025) aim to mitigate this issue by prioritizing high-value data sampling. However, they primarily redistribute resource utilization rather than achieving a genuine reduction in overall resource consumption. We still need to achieve an effective trade-off between training performance and computational cost. To construct the relative advantage, some methods use state-independent baselines to generate advantage values for each action (Williams, 1992; Ahmadian et al., 2024). GRPO  (Shao et al., 2024) and GiGPO (Feng et al., 2025b) compare the rewards of actions or trajectories within groups. In these approaches, the evaluation criterion is derived from the policy itself, which may cause policy optimization to become confined to existing behavior patterns and lead to local optima.
From a human learning perspective, rollout can be seen as repeated practice. Grouping policies resemble trial-and-error learning, where individuals often compare outcomes to a fixed Reference Anchor for more efficient learning. This anchor serves as an objective reference point, distinct from the idealized optimal solutions provided by a critic or the dynamic relative performance within a group, and establishes a more general advantage baseline.
In this paper, we introduce Pre-estimated Value-based Policy Optimization (PVPO), a generalized RL method based on Proximal Policy Optimization (PPO) (Schulman et al., 2017). PVPO adopts a critic-free architecture, is compatible with mainstream group policy RL methods, and maintains low computational cost for grouping, thus effectively combining the strengths of both approaches. Specifically, we use a Reference Model (Ref) to run grouping reasoning and calculate a task-based reward score as an anchor. This anchor serves as the $V$ estimate during RL training, helping to correct the cumulative bias in relative advantage calculations typically observed in large language models (LLMs). In essence, our method decouples $Q$ and $V$ in the grouping policy advantage calculation. The reference anchor is computed in an unsupervised manner and acts as both a supplement and an enhancement to the training dataset, without incurring additional time or memory overhead. In summary, our core contributions are as follows.
  • We propose PVPO, an efficient and generalizable approach to critic-free reinforcement learning. PVPO provides a stable, low-variance, and globally consistent advantage function, effectively mitigating concerns of error accumulation and policy drift during training. As a result, PVPO enables more efficient and robust policy optimization while significantly reducing spatio-temporal overhead.
  • We introduce a group sampling strategy that offline filters data with unstable accuracy rates to construct high-quality batches, thereby enhancing convergence and learning efficiency. Furthermore, for samples with zero accuracy, we leverage a large-scale LLM to generate ground-truth trajectories, facilitating more effective learning from sparse reward signals.
  • PVPO achieves state-of-the-art performance on multi-step retrieval datasets and demonstrates strong generalization on mathematical reasoning benchmarks. Experimental results indicate that PVPO not only enhances multi-hop question answering and tool-use capabilities, but also improves the overall reasoning ability of LLMs.
2 Related Work
2.1 Agentic Reasoning
Leveraging reinforcement learning to drive search represents an important direction in agentic reasoning (Jin et al., 2025; Jiang et al., 2025). Search-o1 (Li et al., 2025b) integrates an agentic search workflow into the reasoning trajectory. This achieves an elegant integration of search and reasoning, sparking a wave of subsequent optimizations (Qian et al., 2025; Wang et al., 2025; Feng et al., 2025a). Moreover, numerous studies on Retrieval-Augmented Generation (RAG) (Li et al., 2025b; Feng et al., 2025c; Hao et al., 2025) have advanced the capabilities of LLM in tool use and information retrieval.
However, existing studies often directly apply algorithms such as GRPO, which are intrinsically ill-suited to the sparse-reward setting of agentic search. These methods depend on dense token-level rewards, necessitating extensive rollouts to achieve stable advantage estimation. Consequently, the quality of the learning signal becomes tightly coupled with the sample size.
Our PVPO framework is tailored for agentic search by decoupling the advantage function (AA=$Q$-$V$), thereby mitigating sample size dependency. While the actual return ($Q$) leverages the sample size, the advantage baseline ($V$) remains independent of both the current and previous policies. This design ensures a stable learning signal even under severe reward sparsity (e.g., QQ=0), obviating the need for extensive rollouts.
2.2 RL for LLMs
Recently, reward and advantage computation has been redefined through dynamic generation and iterative optimization, substantially enhancing the performance of critic-free RL methods. Some methods construct denser feedback signals by increasing the frequency of reward generation (Bensal et al., 2025; Chen et al., 2024), while others improve reward adherence by incorporating additional training phases into the learning process (Dong et al., 2025). These approaches often overlook the compounding hallucinations arising from repeated sampling and error accumulation from iterative policy updates. Each incremental policy change alters the rollout distribution, resulting in advantage estimates targeting a continually shifting objective and potentially steering the policy toward suboptimal local minima. Moreover, these methods depend heavily on costly online sampling procedures.
Another line of research seeks to recover endogenous rewards from the actor model via reverse engineering, a process that has been mathematically substantiated (Li et al., 2025c; Zhao et al., 2025). This approach eliminates the need for additional training and enables adaptation to diverse evaluation preferences through prompt adjustment. However, the quality of the recovered reward is inherently limited by the base model’s capabilities, and consistently guiding reward signals through prompting remains a significant challenge (Zhao et al., 2021; Lu et al., 2022; Liu et al., 2023).
To address these challenges, the research community has investigated various static approaches. The most prominent is offline reinforcement learning, which optimizes policies using fixed datasets (Kumar et al., 2020; Kostrikov et al., 2022). Another notable class comprises Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its variants (Ethayarajh et al., 2024), which reformulate the objective as a direct fit to fixed preference pairs, reducing the reliance on online sampling but constraining generalization. Simpler static methods, such as weighted behavioral cloning (Xu et al., 2022a; b), offer limited expressive power and theoretical guarantees due to their parsimonious advantage estimation.
To balance efficiency and adaptability in policy optimization, our approach integrates a static $V$ with a dynamic $Q$, ensuring stable advantage estimation and low computational overhead while maintaining responsive adaptation to policy updates.
3 Preliminary
In this section, we review the fundamental concepts of policy optimization in RL, with a particular focus on the role of the advantage function and its various estimation methods.
3.1 Proximal Policy Optimization
Actor-critic methods, such as PPO, train a critic network $V_{\phi}(s)$ to provide a low-variance estimate of the state-value function $V^{\pi}(s)$ of state $s$. The state-value function is used to compute the advantage at each time step $t$, typically via Generalized Advantage Estimation (GAE) (Schulman et al., 2015):
$\begin{split}\hat{A}^{\text{GAE}}_{t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l},\quad\delta_{t}=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t})\text{,}\end{split}$ (1)
where $\lambda$ is a hyper-parameter, $\delta_{t}$ is the temporal difference error at time step $t$, $r_{t}$ is the immediate reward received at time step $t$, $\gamma$ is the discount factor. PPO then optimizes a clipped surrogate objective to update the actor network in a stable manner:
$\begin{split}\mathcal{J}^{\text{PPO}}(\theta)={}&\mathbb{E}_{q\sim P(D),o\sim\pi_{\theta_{\text{old}}}(O|q)}\left[\min\left(r_{t}(\theta)\hat{A}^{\text{GAE}}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}^{\text{GAE}}_{t}\right)\right]\text{,}\end{split}$ (2)
where $q$ are questions sampled from the dataset $D$, $o$ are outputs sampled from the old policy $\pi_{\text{old}}$, importance sampling ratio $r_{t}(\theta)=\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}$, $\epsilon$ is the clipping range of $r_{t}(\theta)$.
3.2 Group Relative Policy Optimization
Since the critic network is typically as large as the actor network, it adds substantial memory and computational burden. Critic-free methods, such as GRPO, eliminate this costly component by estimating the advantage directly from rewards.
For each question, GRPO generates a group of outputs $\{o_{i}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. The advantage for each output $o_{i}$ is then calculated based on normalized reward $\mathbf{r}$ relative to the group:
$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}\text{.}$ (3)
This critic-free advantage estimate is then used to optimize a PPO-like objective function:
$\begin{split}&\mathcal{J}^{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(D),\{o_{i}\}\sim\pi_{\theta_{\text{old}}}(O|q)}\\ &\quad\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\biggl{\{}\min\left(r_{i,t}(\theta)\hat{A}_{i,t},\text{clip}\left(r_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right)-\beta D_{KL}[\pi_{\theta}||\pi_{\text{ref}}]\biggr{\}}\Biggr{]}\text{,}\end{split}$ (4)
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$, $D_{KL}$ is the KL divergence between the trained policy $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$, $\beta$ is a hyper-parameter.
Figure 1: The architecture of PVPO. Reference model updates $R^{\text{ref}}$ at fixed steps, maintaining value stability and improving the performance lower bound. Reward manager do not restrict the generation of reward.
4 Methodology
In this section, we will introduce our efficient and effective RL algorithm PVPO. The architecture is illustrated in Figure 1. PVPO optimizes the policy via the following objective:
$\begin{split}&\mathcal{J}^{\text{PVPO}}(\theta)=\mathbb{E}_{q\sim P(D),\{o_{i}\}\sim\pi_{\theta_{\text{old}}}(O|q)}\\ &\quad\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\biggl{\{}\min\left(r_{i,t}(\theta)\hat{A}^{\text{PVPO}}_{i,t},\text{clip}\left(r_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}^{\text{PVPO}}_{i,t}\right)-\beta D_{KL}[\pi_{\theta}||\pi_{\text{ref}}]\biggr{\}}\Biggr{]}\text{.}\end{split}$ (5)
where
$r_{i,t}(\theta)=\begin{cases}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},&\text{if }o_{i}\notin\text{GT Traj}.\\ \frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{gt}}}(o_{i,t}|q,o_{i,<t})},&\text{if }o_{i}\in\text{GT Traj}.\end{cases}$ (6)
4.1 Static V Estimate
In actual policy optimization, the current method is to operate at the group level rather than through single sampling. For problem $q$, we use the current policy $\pi_{\theta}$ to generate $N$ independent trajectories $\mathcal{T}=\{\tau_{1},\tau_{2},...,\tau_{N}\}$ and obtain the corresponding rewards $\mathbf{r}=\{R(\tau_{1}),R(\tau_{2}),...,R(\tau_{N})\}=\{r_{1},r_{2},...,r_{N}\}$. For any step $(s_{i,t},a_{i,t})$ in a specific trajectory $\tau_{i}$, the unbiased Monte Carlo estimate of the action value $Q^{\pi}(s_{i,t},a_{i,t})$ is the final reward $r_{i}$ observed in that trajectory. We refer to this as the Dynamic Q Estimate because it directly reflects the result of a single rollout of the current policy:
$\hat{Q}_{\text{dyn}}(\tau_{i})=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau_{i})]=r_{i}\text{.}$ (7)
Considering that reward $r_{i}$ is given after the generation of trajectory $\tau_{i}$, the trajectory generation process is regarded as atomic actions $a_{i}=\tau_{i}$ executed from $s_{i,0}$. This atomicity makes the reward distribution of the intermediate state $s_{i,t}$ only depend on initial state $s_{i,0}$ ($s_{0}$) and $\pi_{i}$. Consequently, the expected return of the policy is equal to the state value of the initial state $V^{\pi}(s_{0})$. A natural estimation method is to approximate this expectation using the empirical mean of all rewards in the current group. This is the approach adopted by on-policy methods such as GRPO, which we refer to as Dynamic V Estimate:
$\hat{V}_{\text{dyn}}(s_{0})=\hat{V}_{\text{dyn}}(\mathcal{T})=\frac{1}{N}\sum_{j=1}^{N}r_{j}=\text{mean}(\mathbf{r})\text{.}$ (8)
So we obtain the sparse advantage estimate for trajectory $\tau_{i}$ in the on-policy method:
$\begin{split}\hat{A}_{\text{dyn}}(\tau_{i},s_{0})&=\hat{Q}_{\text{dyn}}(\tau_{i})-\hat{V}_{\text{dyn}}(s_{0})\\ &=r_{i}-\text{mean}(\mathbf{r})\text{.}\end{split}$ (9)
This formula clearly shows that the advantage is calculated as the difference between the immediate reward and the average performance of the current policy $\pi_{\theta}$ within the group. However, $\hat{V}_{\text{dyn}}$ fluctuates wildly with each sampling of the group and is directly affected by $\pi_{\theta}$, introducing significant instability, especially when the group size is not large enough. To more effectively mitigate the instability associated with dynamic $V$ estimation, we propose substituting it with a more robust fixed $V$ estimate.
The ideal baseline should represent a Reference Anchor that does not change with current policy iterations. Therefore, we use the expected return of a fixed reference policy $\pi_{ref}$ (e.g., the initial policy model) as our Static V Estimate $\hat{V}_{\text{sta}}$. The baseline can be accurately estimated in advance by sampling the reference policy $\pi_{ref}$ M times, and update at fixed steps during training process:
$\hat{V}_{\text{sta}}(s_{0})=\frac{1}{M}\sum_{j=1}^{M}r^{\text{ref}}_{j}=\text{mean}(\mathbf{r^{\text{ref}}})\text{.}$ (10)
This stable static baseline replaces the unstable dynamic baseline in formula 8. We finally obtain the advantage function of PVPO, which is applicable to all sparse reward fields:
$\begin{split}\hat{A}^{\text{PVPO}}(\tau_{i},s_{0})&=\hat{Q}_{\text{dyn}}(\tau_{i})-\hat{V}_{\text{sta}}(s_{0})\\ &=r_{i}-\text{mean}(\mathbf{r^{\text{ref}}})\text{.}\end{split}$ (11)
For any step $(s_{i,t},a_{i,t})$ of any trajectory $\tau_{i}$, the advantage can also be written as follows:
$\hat{A}^{\text{PVPO}}_{i,t}$ $=r_{i}-\text{mean}(\mathbf{r}^{\text{ref}})\text{,}$ (12)
$r_{i}$ $=\begin{cases}\max(0.1,\,r_{\text{acc}}),&\text{if format is correct.}\\ 0,&\text{if format is incorrect.}\end{cases}$ (13)
$r_{\text{acc}}$ $=\begin{cases}\text{F}_{1}(a_{\text{pred}},\,a_{\text{gt}}),&\text{if }L_{\text{pred}}\geq n\cdot L_{\text{gt}}.\\ \text{CEM}(a_{\text{pred}},\,a_{\text{gt}}),&\text{if }L_{\text{pred}}<n\cdot L_{\text{gt}}.\end{cases}$ (14)
where $a_{*}$ denotes the answer, $L_{*}$ denotes the length of the answer, $_{pred}$ denotes the prediction, and $_{gt}$ denotes the ground truth. $r^{ref}_{i}$ is similar. F1 denotes the standard word-level F1 score and CEM denotes Cover Exact Match. $n$ represents the multiple of the text length, which we set to $3$ by default. This reward is used for agentic search, following Hao et al. (2025).
In summary, $\hat{Q}_{\text{dyn}}(\tau_{i})$ is obtained from the immediate reward of on-policy $\pi_{\theta}$ rollout. It reflects the current performance of the policy and is highly adaptive. The Static V Estimate $\hat{V}_{\text{sta}}(s_{0})$ is obtained from the average reward of the reference policy $\pi_{\text{ref}}$ pre-rollout. It provides a stable and low-variance performance baseline.
4.2 Group Sampling
Inspired by DAPO’s dynamic sampling strategy, we also assess the accuracy of sample rollouts while continuing to utilize the reference model for offline rollouts. For each sample, the mean accuracy of the rollouts serves as the filtering criterion.
Specifically, samples are categorized into three groups:
  • Samples with a mean accuracy of 1 are excluded from the training set, as they are considered too trivial to facilitate effective learning.
  • Samples with a mean accuracy strictly between 0 and 1 are retained, given their nonzero advantage.
  • For samples exhibiting a mean accuracy of 0, an additional rollout is conducted using a larger LLM for further evaluation.
The larger LLM can correctly answer some of these samples. We cache these Ground Truth Trajectories (GT Traj) and their probability distributions. During policy training, a GT Traj is injected by replacing one of the generated rollouts for these specific samples. This method mitigates the sparse reward issue commonly encountered with complex samples. In the absence of guidance, the LLM may fail to obtain any positive feedback through unguided exploration. By providing a reference trajectory, the model receives an explicit demonstration, which jumpstarts learning by offering a clear example of a successful reasoning process.
5 Expriments Setting
5.1 Metrics
For multi-hop question-answering tasks, we employ answer accuracy (Acc, %) and LLM-as-a-Judge (LasJ, %) (Song et al., 2025) as benchmarks. For mathematical reasoning tasks, we employ answer accuracy (Acc, %) as the evaluation metric, specifically reporting the mean accuracy across 32 independent rollouts for each sample (mean acc@32).
5.2 Data Collection and Processing
For multi-hop question-answering tasks, we conducted extensive experiments on four multi-step retrieval datasets, including Musique (Trivedi et al., 2022), 2WikiMultiHopQA (2Wiki) (Ho et al., 2020), HotpotQA (Yang et al., 2018), and Bamboogle (Bam) (Press et al., 2023). We trained on Musique, with the original dataset size of 20k and a filtered size of 8k after Group Sampling. Evaluations were done on the full development or test sets of these datasets.
For mathematical reasoning tasks, we trained on DAPO-Math-17k-Processed (Yu et al., 2025), with an original size of 17k and a filtered size of 10k after Group Sampling. We evaluated on five test sets: DAPO-AIME-2024 (AI-MO, 2024; Bytedance & Tsinghua-SIA, 2025), AIME-2025 (Lin, 2025), MATH500 (Lightman et al., 2024; HuggingFaceH4, 2023), AMC23 (AI-MO, 2024), and Olympiad (He et al., 2024).
5.3 Environment and Hyperparameters
All experiments ran on a server with an Intel(R) Xeon(R) Platinum 8369B CPU and eight NVIDIA A100-SXM4-80GB GPUs. We used Qwen2.5-7B-Instruct as the baseline model and Qwen2.5-72B-Instruct as the large LLM to generate GT Traj. $R^{\text{ref}}$ was updated every 500 steps. The number of pre-samples $M$ in the reference model was set equal to the number of samples $N$ in the policy model.
Our multi-hop question answering task used the ReSearch (Chen et al., 2025) framework, with batch size 8, rollout count $N=5$, and 1,000 training steps. DynaSearcher added “kg_filter” during inference, but we used direct inference throughout.
The mathematical reasoning task used the veRL (Sheng et al., 2025) framework, with batch size 32, rollout count $N=16$, and 1,000 training steps. GRPO’s “loss_agg_mode” was set to the default “token-mean” instead of “seq-mean-token-mean”. For evaluation, we set “val_temperature” = 0.6 and “val_top_p” = 0.95.
6 Experiments
In this section, we run a series of experiments. First, we test our method on multi-hop question answering to validate its effectiveness in the agent domain. Then, we test it on mathematical reasoning to show its generalizability. These experiments help us answer the following questions:
Q1: Compared to existing RL methods, can PVPO achieve better performance?
Q2: Does PVPO show generalizability across different domains?
Q3: Can PVPO reduce the temporal resource consumption required for training?
Q4: Does PVPO have better stability than baseline methods during training?
Q5: Does training with Static V Estimate really help the model?
6.1 Main Performance Comparison
To answer Q1, we evaluated PVPO on multi-step retrieval datasets and compared the results with two types of benchmarks. The results are shown in Table 1.
Table 1: Performance comparisons between PVPO and the baselines on multi-step retrieval datasets. The best and second best results are bold and underlined, respectively. $\dagger$$\ddagger$} represents in-domain/out-of-domain datasets.
Musique 2Wiki HotpotQA Bamboogle Avg Acc
Acc LasJ Acc LasJ Acc LasJ Acc LasJ Acc LasJ
Prompt Based
Qwen2.5-7B-Instruct 5.10 13.48 27.90 29.30 22.40 31.04 12.80 17.12 17.05 22.74
DeepSeek-R1 32.00 40.70 57.50 59.40 43.00 58.30 66.40 76.64 49.73 58.76
O4-mini 38.00 44.10 61.50 67.40 49.50 67.40 74.40 84.16 55.85 65.77
GPT-4.1-global 31.00 40.90 58.00 58.50 44.50 57.70 51.20 61.60 46.18 54.68
Gemini-2.5-pro 42.50 50.80 70.00 71.20 53.00 71.10 75.20 84.48 60.18 69.40
Train Based
Qwen2.5-7B-Instruct
ReSearch 13.40 24.90 41.70 43.26 33.40 47.26 36.00 43.20 31.13 39.66
GRPO on ReSearch 33.40 46.72 60.80 67.02 54.50 63.68 45.60 54.40 48.58 57.96
GRPO on DynaSearcher 38.90 52.04 74.30 76.77 62.70 68.32 51.20 58.72 56.78 63.96
PVPO on ReSearch 36.50 51.44 70.10 72.36 65.50 72.34 45.60 54.32 54.43 62.62
PVPO on DynaSearcher 46.90 59.44 77.70 80.62 69.00 78.44 50.40 59.68 61.00 69.55
6.1.1 Performance of LLMs
We compared PVPO with the zero-shot inference performance of leading open-source and closed-source LLMs, including GPT-4.1-0414-global, DeepSeek-R1-0528, O4-mini-0416-global, and Gemini-2.5-pro-0325. The results show that the 7B model trained with PVPO improved significantly, reaching 3.6 times the accuracy of the original model. It outperforms all general LLMs on in-domain datasets. On the out-of-domain dataset, its performance is not as high as in zero-shot settings. This may be due to differences in training data and model capacity, since these leading models might have seen similar data during pre-training or fine-tuning, while the 7B model has limited capacity. However, its average accuracy is 8 percentage points higher than the average of the four LLMs and about 1 percentage point higher than the best-performing LLM.
6.1.2 Performance of Agentic Retrieval Methods
To enable a more fair comparison, we also compared PVPO with agentic retrieval methods that require training. In this experiment, all methods were trained for 1,000 steps. We included the direct inference methods of ReSearch and DynaSearcher, as well as a baseline that trained these two methods using RL with GRPO. We observed that the performance of PVPO improved by more than 5 percentage points on average compared to GRPO. The result demonstrates the consistent superiority of PVPO across agentic retrieval methods.
Overall, this answers Q1: Compared to existing reinforcement learning methods, PVPO achieves state-of-the-art performance.
Table 2: Performance comparisons between PVPO and GRPO under different scale models on mathematical reasoning datasets.
MATH500 AMC23 Olympiad AIME-2024 AIME-2025 Avg Acc
Qwen2.5-7B-Instruct
GRPO 78.66 49.20 42.26 14.26 12.83 39.44
PVPO 80.30 52.02 44.62 14.86 14.70 41.30
Qwen2.5-14B-Instruct
GRPO 82.22 55.94 49.70 17.14 16.90 44.38
PVPO 83.64 56.78 50.72 19.24 17.74 45.62
6.2 Generalization Evaluation
We shifted from multi-hop question answering to single-shot mathematical reasoning. We chose several math reasoning datasets with varying difficulty, from basic arithmetic to challenging olympiad-level problems. The goal is to test general reasoning ability in different fields and situations. In the experiment, we compared PVPO with GRPO. The results are shown in Table 2. In terms of average performance, PVPO outperforms GRPO at both model scales. On the 7B model, PVPO scores 1.89 percentage points higher than GRPO. When scaled up to 14B, PVPO continues to lead. PVPO shows clear advantages on most datasets.
Overall, PVPO achieved the best results on all five math reasoning datasets. This answers Q2: PVPO is generalizable and keeps stable performance across different fields and tasks.
Figure 2: Training efficiency study of PVPO on mathematical reasoning datasets.
6.3 Training Efficiency Analysis
6.3.1 Convergence Speed
We compare PVPO and GRPO to show PVPO’s faster convergence speed. During training, we tracked accuracy and rewards, as shown in Figure 2. Figure 2 (a) shows that after about 500 steps, PVPO reaches the accuracy level that GRPO gets after 1,000 steps. Figure 2 (b) shows that PVPO starts with a higher reward than GRPO and grows at a faster rate. Group Sampling provides high-gain samples during training, and the static advantage baseline gives a clear direction for policy optimization. Overall, PVPO reaches the target accuracy with fewer steps and converges faster.
Figure 3: Group Sampling study on datasets from different fields. The Acc is the mean of the answer accuracies from M trajectories rolled out by the reference model. M=5 in Figure (a) and M=16 in Figure (b).
6.3.2 Total Time Consumption
[…]
Figure 4: Training stability study of PVPO on multi-step retrieval datasets.
6.4 Stability Evaluation
We tracked PVPO training metrics to show its stability. Figure 4 (a) shows PVPO achieves a much higher average reward than GRPO. With similar KL divergence in Figure 4 (b), this improvement comes not from more aggressive updates, but from better gradient direction estimates. As shown in Figure 4 (c), PVPO has much lower advantage variance, leading to more reliable and consistent update directions. PVPO also maintains exploration without losing stability. Figure 4 (d) shows it keeps higher policy entropy under a similar KL constraint, which helps avoid premature convergence to a local optimum. Overall, PVPO addresses key problems in RL: it supports high exploration, low variance, and high rewards. This evidence answers Q4: PVPO achieves more stable training than existing methods.
6.5 Case Study: Low Sampling Budget
We ran a case study to address Q5. We reduced the number of rollouts from 5 (used in the main experiments) to 2. For comparison, we report GRPO’s performance with a full budget. Figure 5 (a) shows that PVPO with a low budget remains close to the fully budgeted GRPO. We calculate computational cost by multiplying the number of rollouts with the average number of tool calls in trajectories. As shown in Figure 5 (b), PVPO’s average cost is only 4.3, which is much lower than GRPO’s 11.7. PVPO achieves 97% of GRPO’s performance (55.0% vs. 56.8%) while using less than 40% of the computational cost. This strong sample efficiency comes from the high-quality, low-variance training signals provided by Static V Estimate. The model can update its policy efficiently using fewer rollouts. This answers Q5: model training clearly benefits from Static V Estimate.
Figure 5: Low sampling budget study of PVPO on multi-step retrieval datasets. The NN denotes the number of trajectories in each single rollout. NN=5 is the full budget and NN=2 is the low budget.
7 Conclusions
In this paper, we propose PVPO, an efficient critic-free reinforcement learning algorithm designed to optimize policy learning for complex tasks. By introducing a Static V Estimate as an external advantage reference and integrating it with group sampling for effective data filtering, PVPO addresses the limitations of extensive sampling and biased intra-group comparisons inherent in prior methods. Our approach yields stable, low-variance training signals, accelerates convergence, and significantly reduces computational costs. Extensive experiments across nine diverse benchmarks in multi-hop question answering and mathematical reasoning demonstrate that PVPO achieves state-of-the-art performance and strong generalization, even with small-scale models and limited resources. PVPO introduces substantial improvements in reasoning and tool use, supports scalable training, and ensures consistent performance, thereby demonstrating strong potential for widespread real-world application.